# A simple, flexible method for timing cross-calibration of space missions

### v1

@@ 2.3 Need for time system consistency
Whether the barycentering is done before or after folding, the procedure involves a number of very delicate steps and decisions, which can introduce systematic errors in the TOAs. The solar system barycenter is defined as the center of mass of the solar system, which is not at the center of the Sun, but rather at a point that moves around it due to the gravitational pull of the planets and the other objects in the solar system. This is done through the use of solar system ephemerides, which are mathematical models that describe the positions and motions of the solar system bodies. In astronomy it is customary to use the JPL development ephemerides, which are produced by the Jet Propulsion Laboratory and are the result of the numerical integration of the equations of motion of solar system objects.
@@
The JPL ephemerides are updated regularly and are indicated by a number, such as DE200, DE405, DE421, DE430, etc. The choice of the ephemeris can have a significant impact on the barycentering procedure, introducing systematic errors in the TOAs by up to a few milliseconds, which is often a significant fraction of the pulsar’s spin period.
Another issue is the choice of the definition of the barycentric time, which can be for example the barycentric dynamical time (TDB) or the barycentric coordinate time (TCB). The two systems are very similar, but they differ in the way they account for relativistic effects.
@@
The effect is macroscopic and it is often used to refine the pulsar’s position in the sky with a precision comparable to high-resolution VLBI or optical observations. Proper motion can also produce a significant effect on the TOAs and be measured through pulsar timing. One has to pay attention to the choice of coordinates: the right ascension and declination are usually given in the J2000 equinox in the ICRS coordinate frame, but DE200 uses the older FK5 reference frame, which differs by a fraction of an arcsecond from the ICRS frame and produces a shift up to $50\,\rm\mu s$ in our processing.
Lower-order effects, on the order of micro seconds, can arise from the choice of clock corrections to be applied to the barycentric time.
@@
2.4 Using the Jodrell Bank ephemeris
@@
The Jodrell Bank monthly ephemerides (hereafter JBE; A. G. Lyne et al. 1993) tracks the rotational history of the Crab pulsar over time, publishing a new timing solution roughly every month. This priceless dataset, with a time span of almost four decades, is the ideal reference for the timing of the Crab pulsar. Data are shared in two different formats: a space-separated text file with the spin frequency, one spin derivative and a TOA at the barycenter of the solar system, and a CGRO format [^3] https://www.jb.man.ac.uk/research/pulsar/crab/CGRO_format.html file with two spin derivatives and a TOA at the geocenter. Say our X-ray observation was performed at MJD 56000. The corresponding line in the CGRO-format file gives the solution between MJDs 55987–56018, comprising of a spin frequency $\dot{\nu}$=29.7025600117591 Hz, a spin derivative $\dot{\nu}=3.70689\mathrm{e}^{-10}$ Hz s-1 , and a second derivative $\ddot{\nu}=1.95\mathrm{e}^{-20}$ Hz s-2. Moreover, this solution includes a TOA at the geocenter (MJD 56003.000000285). These ingredients should be enough for a model describing the phase of the Crab during the month with an rms scatter of 0.6 milliperiods, also declared in this solution. However, the JBE ephemerides publish the timing solution at the solar system barycenter using the DE200 FK5 reference frame, which is quite outdated, and not accounting for proper motion, so assuming that the pulsar’s position is fixed in the sky. The reason is consistency: users of the ephemerides do not need to worry about using different reference frames for datasets spanning different epochs (Lyne, priv. comm.). The use of the JBE for X-ray timing studies, therefore, forces to barycenter the data using the same set of outdated ephemerides and position, losing flexibility when gathering data from all different missions (that might have been barycentered using different conventions).
3 Supporting multiple JPL ephemerides at once
@@
    we fit a new timing solution to these TOAs, using the DE430 ephemeris and the same position used for barycentering the X-ray data;
@@
    we verify that the scatter of the TOAs with the new solution is much smaller than the declared rms scatter of the JBE (typically in the order of the noise injected in the random TOAs)
@@
At this point, we have a new timing solution that is consistent with the JBE, but using the DE430 ephemeris. This procedure can easily be done using the PINT (J. Luo et al., 2021) software package, that supports a large set of timing models and JPL ephemerides. The solution can be saved in a parameter file in PINT (J. Luo et al., 2021) format, containing all the relevant information. The geocenter arrival time from the JBE is encoded in the new model as TZRMJD, using 0 as observatory. Special care must be put in the choice of coordinates relevant to a given ephemeris. DE200 uses the FK5 frame, whilst DE4xx uses ICRS. The difference of arrival times using the wrong coordinate frames can be substantial, in the order of milliseconds.
3.1 Folding
@@
Once we have obtained a suitable timing solution for our observation, we can use it to fold the event arrival times and obtain a pulse profile. The pulse phase is calculated with the usual Taylor expansion in Eq. 1, using the timing model obtained in Section 3. and the pulse profile is just a histogram of pulse phases, between 0 and 1, using $N$ phase bins. The frequency and frequency derivatives are encoded in the timing solution. At this point, the fundamental missing ingredient is $\phi_{0}$. This has to be calculated so that the reference TOA in the JBE solution is at phase 0. If we do this, since the phases are already referred to the radio TOA, our folded profiles will have automatically the phase 0 at the peak of the radio TOA. All the relevant time conversions are done by PINT, by using the TEMPO2-format parameter file created with the procedure in Section 3. The phase of each photon can be calculated by properly converting the event arrival times from mission-elapsed seconds to MJD and using the TimingModel.get_phase method. Observations spanning multiple JBE solutions (e.g. from Fermi, IXPE) were split in the relevant time ranges. In very high-throughput observations with more than 100M photons (e.g. frequently from NICER) we calculated a TOA every $\sim$50M photons.
3.2 Deadtime correction of pulse profiles
@@
Figure 1: Example of deadtime correction of a NuSTAR pulse profile (ObsID 10002001009), using the method described in Section 3.2. The blue line shows the exposure correction, the red line the raw profile and the black line the deadtime-corrected profile.
@@
For missions affected by dead time, like NuSTAR, we corrected the pulse profiles using the PRIOR column in the event list, which gives the live time before the current event, largely following the procedure used by K. K. Madsen et al. (2015). In short, the live time before each event was used to estimate the effective exposure in each bin of the pulse profile. For example, let us consider photon arrived at time $t_{0}$, with a value in the PRIOR column of 1 ms. This means that the live time before this photon started at $t_{p}=t_{0}-1\,\mathrm{ms}$. Let us call $\phi_{0}$ the phase of the pulse profile corresponding to $t_{0}$, and $\phi_{p}$ the phase corresponding to $t_{p}$. Our procedure will add 1 to all phase bins between $\phi_{p}$ and $\phi_{0}$, properly accounting for fractional bin coverage and phase wraps at the start and end of each pulse profile. In the end, the exposure will be normalized to a maximum of 1, and the pulse profile divided by this effective exposure (so that more deadtime-affected pulse bins will be amplified). This procedure is made available in the code pulse_deadtime_fix (M. Bachetti, 2025a). The result of this procedure is shown in Figure 1.
@@
between 12 and 24 counts per second to each bin of the folded profile, following D. Pease & H. Donnelly (1998).
3.3 Residual calculation
@@
Figure 2: Folded profile from NICER (obsid 101301013) modeled with the combination of symmetric and asymmetric Lorentzians described in Section $3.$
@@
Once we have the folded profile, as we saw in Section 3.1, the phase 0 of the profile should coincide with the main radio peak. As very well known from the literature (A. H. Rots et al., 2004), the peak in the X-rays will not coincide with the radio peak, and we need a way to measure this delay, that can then be used to compare the performance of different high-energy missions.
@@
There are mainly two methods used in the literature to calculate the phase of a pulsar: the alignment of the folded profile with a template (e.g. FFTFIT, J. H. Taylor 1992) or fitting a model to the pulse profile (e.g. J. Nelson et al., 1970). The first method is the most stable in ideal conditions, when the template for the pulse is correct, bringing precise measurements even in noisy data. However, if the template is not correct, as can happen if the response of the instrument used to calculate the template is different from the instrument used to obtain the folded data, or worse, if the folding is not perfect due to issues with the timing systems, this method can give unrealistically small errors and/or align to the wrong peaks. In the case of the Crab pulsar, where low signal-to-noise is seldom a problem, the second method yields meaningful results both in “good” datasets without requiring an exact template, and in datasets affected by instrumental issues, where the template adapts to an imperfect profile measuring the peak position and providing a reasonable error bar.
@@
For this work, we chose to model each peak as a combination of a symmetric Lorentzian curve plus an asymmetric Lorentzian curve. The phases of the two Lorentzians composing each peak are linked (not more than a difference of 0.02 in phase). The folded profile changes considerably at different energies, until the secondary peak becomes comparable to the first peak making the identification less obvious. However, the secondary peak maintains a distance of $\sim 0.4$ in phase from the primary, so that the two peaks can easily be identified by their phase distance. The error on the arrival time can be calculated from the error of the centroid of the sharpest Lorentzian. We verified that the scatter of TOA residuals is indeed comparable to the error bars estimated in this way. The fit quality can be seen in Figure 2.
3.4 Validation
@@
Figure 3: Results of the validation procedure for the model refitting described in Section 3.4. We fit a DE430 model to DE200-simulated TOAs, then simulated TOAs with the new model and calculated the residuals to the original DE200 model. For each instance of the simulation, we plot the mean residual in the upper panel and the rms of the residuals in the lower panel. The procedure works remarkably well, with residuals and rms well below the quoted accuracy of the radio ephemeris.
@@
Figure 4: End-to-end validation of the procedure in Section 3 using NICER datasets. (Left) Difference between the TOAs calculated from datasets barycentered using the DE200, DE430 and DE440 JPL ephemerides. (Left) a histogram of the corresponding cases. We did not consider observations where the TOA uncertainty was $15\,\rm\mu s$. This demonstrates that the method described in this paper is able to calculate TOAs from a variety of JPL ephemerides and positions, and that the differences are typically $5\,\rm\mu s$, with a possible systematic shift by $\sim 2\rm\mu s$ that we are investigating. The uncertainty of these TOAs is $\sim 5\,\rm\mu s$.
@@
Starting from a model of the Crab pulsar defined in the DE200 ephemeris, we simulated one month of geocentered TOAs as in Section 3, we fit a DE430 model to them, simulated geocentered TOAs with the new model, and calculated the residuals to the initial model. Results are shown in Figure 3. We found that, for the method to work with a precision of more than 5 $\rm\mu s$, quadruple precision floating points have to be available in the computer doing the analysis. For example, in recent Silicon-based Mac machines, the scatter in Figure 3 increased by a factor $\sim$20, with strong patterns dependent on the pulse frequency and the day of year. The problem did not appear on Intel-based Macs, Linux machines, or when using the Rosetta emulator on Silicon Macs.
Difference between ephemerides in barycentering
We compared the phases of photons from NuSTAR observations of the Crab pulsar, barycentered using the DE200 and DE430 ephemerides and folded using the timing models obtained with our method. Differences were below a few $\rm\mu s$(Figure 4).
Calculation of local frequency
We also calculated the local frequency of the pulsar by applying the quasi-fast folding algorithm (M. Bachetti et al., 2020), as implemented in Stingray (D. Huppenkothen et al., 2019), to each observation, searching a small frequency-frequency derivative plane around the radio solution.
3.5 Pipeline

### v3

@@ II.3 Consistency in ephemerides, coordinates, reference frames
Whether the barycentering is done before or after folding, the procedure involves a number of very delicate steps and decisions, which can introduce systematic errors in the TOAs. The solar system barycenter is defined as the center of mass of the solar system, which is not at the center of the Sun, but rather at a point that moves around it due to the gravitational pull of the planets and the other objects in the solar system. This is done through the use of solar system ephemerides, which are mathematical models that describe the positions and motions of the solar system bodies. In X-ray astronomy it is customary to use the JPL development ephemerides, which are produced by the Jet Propulsion Laboratory and are the result of the numerical integration of the equations of motion of solar system objects [^5] Alternative ephemerides exist, like the ones from the Institute of Celestial Mechanics and Ephemeris Calculation (IMCCE), called INPOP, and the ones from the Institute of Applied Astronomy, Russian Academy of Sciences (IAA RAS), called EPM, providing similar precision (Moiseev and Emelyanov, 2024)
@@
The JPL ephemerides are updated regularly and are indicated by a number, such as DE200, DE405, DE421, DE430, etc. The choice of the ephemeris can have a significant impact on the barycentering procedure, introducing systematic errors in the TOAs by up to a few milliseconds, which is often a significant fraction of the pulsar’s spin period. Another issue is the choice of TDB or TCB, as mentioned in Appendix A.
@@
The effect is macroscopic and it is often used to refine the pulsar’s position in the sky with a precision comparable to high-resolution VLBI or optical observations. Proper motion can also produce a significant effect on the TOAs and be measured through pulsar timing. One has to pay attention to the choice of coordinates: the right ascension and declination are usually given in the J2000 equinox in the ICRS coordinate frame [^6] More appropriately, ICRS is the name of the coordinate system, while the frame should be called ICRF; however, this rather confusing naming is used by barycorr and we decided to use the same for simplicity., but JPL ephemerides DE200-DE402 use the older FK5 reference frame, which differs by a fraction of an arcsecond from the ICRS frame and produces a shift up to $50\,\rm\mu s$ in our processing.
@@
II.4 Using the Jodrell Bank ephemeris
@@
The Jodrell Bank monthly ephemerides (hereafter JBE; Lyne et al. 1993) tracks the rotational history of the Crab pulsar over time, publishing a new timing solution roughly every month. This priceless dataset, with a time span of almost four decades, is the ideal reference for the timing of the Crab pulsar. Data are shared in two different formats: a space-separated text file with the spin frequency, one spin derivative and a TOA at the barycenter of the solar system, and a CGRO format [^7] https://www.jb.man.ac.uk/research/pulsar/crab/CGRO_format.html; note that the coordinates quoted in this file correspond to (R.A., Dec.) 05:34:31.972 +22:00:52.070, 4.5 milliarcseconds from the official JBE coordinates 05:34:31.97232 +22:00:52.069, and that this solution does not quote a dispersion measure, so it is only useful for non-radio data file with two spin derivatives and a TOA at the geocenter, defined as the peak of the radio profile at infinite frequency (i.e., corrected for interstellar dispersion). Say our X-ray observation was performed at MJD 56000. The corresponding line in the CGRO-format file gives the solution between MJDs 55987–56018, comprising of a spin frequency $\dot{\nu}$=29.7025600117591 Hz, a spin derivative $\dot{\nu}=3.70689\times 10^{-10}$ Hz s-1 , and a second derivative $\ddot{\nu}=1.95\times 10^{-20}$ Hz s-2. Moreover, this solution includes a TOA at the geocenter (MJD 56003.000000285). These ingredients should be enough for a model describing the phase of the Crab during the month with an rms scatter of 0.6 milliperiods, also declared in this solution. However, the JBE ephemerides publish the timing solution at the solar system barycenter using the DE200 FK5 reference frame, which is quite outdated, and not accounting for proper motion, so assuming that the pulsar’s position is fixed in the sky to the J2000 position (R.A., Dec.) 05:34:31.97232 +22:00:52.0690 in the FK5 frame, or (R.A., Dec.) 05:34:31.970490 +22:00:52.087768 in the ICRS frame [^8] Throughout the paper, we use the functions in astropy.coordinates to perform the coordinate conversion, e.g. defining coordFK5 = SkyCoord(ra, dec, frame="fk5") and then using coordFK5.icrs for the conversion to ICRS. This, despite the significant proper motion by $\sim 13$ milliarcsec/yr, which implies that the pulsar has moved by more than 0.5″from the position quoted in the JBE (Ng and Romani, 2006; Kaplan et al., 2008; Gaia Collaboration, 2020; Lin et al., 2023). A parallax of $0.53\pm 0.06$ mas has also been measured (Lin et al., 2023), but its effect on the TOAs is negligible ($\lesssim 1\,\rm\mu s$) for the scope of this paper. See Table 1 for a comparison of the Crab pulsar coordinates quoted by the JBE and Gaia Collaboration (2020) in different reference frames.
Table 1: Crab pulsar coordinates from JBE and Gaia (Gaia Collaboration, 2020) expressed in different reference frames with precision truncated to three digits. The uncertainties are much smaller than the differences between the measurements, which are dominated by the systematic differences between the reference frames.
Reference frame Jodrell Bank (Epoch 1970.0999The Crab notes document does not state an epoch. A review by Taylor et al. (1993) quotes the same position citing McNamara (1971), ref. 156 in Table 1. In McNamara, the epoch is 1971.0 but the right ascension is different –05:31:31.428 (B1950) instead of 05:31:31.405. The position is assumed to be fixed in the JBE, in any case, so the epoch has no practical effect.   ) Gaia (Epoch 2016.0)
RA Dec RA Dec
FK4 (B1950) 05:31:31.428 +21:58:54.402 05:31:31.405 +21:58:54.466
FK5 (J2000) 05:34:31.972 +22:00:52.069 05:34:31.949 +22:00:52.135
ICRS (J2000) 05:34:31.970 +22:00:52.088 05:34:31.947 +22:00:52.154
The reason is consistency: users of the ephemerides do not need to worry about using different reference frames for datasets spanning different epochs (Lyne, priv. comm.). The use of the JBE for X-ray timing studies, therefore, forces to barycenter the data using the same set of outdated ephemerides and position, losing flexibility when gathering data from all different missions (that might have been barycentered using different conventions).
III Supporting multiple JPL ephemerides at once
@@
    We fit a new timing solution to these TOAs, using the DE430 ephemeris and the same ICRS position used for barycentering the X-ray data;
@@
[…]
@@
At this point, we have a new timing solution that is consistent with the JBE, but using the DE430 ephemeris. This procedure can be implemented straightforwardly using the PINT (Luo et al., 2021) software package, that supports a large set of timing models and JPL ephemerides. The solution can be saved in a parameter file in PINT (Luo et al., 2021) format, containing all the relevant information. The geocenter arrival time from the JBE is encoded in the new model as TZRMJD, using 0 as observatory. Special care must be put in the choice of coordinates relevant to a given ephemeris. DE200 uses the FK5 frame, whilst DE402 and later uses ICRS. The difference of arrival times using the wrong coordinate frames can be substantial, in the order of milliseconds.
III.1 Folding
@@
Once we have obtained a suitable timing solution for our observation, we can use it to fold the event arrival times and obtain a pulse profile. The pulse phase is calculated with the usual Taylor expansion in Eq. 1, using the timing model obtained in Section III, and the pulse profile is just a histogram of pulse phases, between 0 and 1, using $N$ phase bins. The frequency and frequency derivatives are encoded in the timing solution. At this point, the fundamental missing ingredient is $\phi_{0}$. This has to be calculated so that the reference TOA in the JBE solution is at phase 0. If we do this, since the phases are already referred to the radio TOA, our folded profiles will have automatically the phase 0 at the peak of the radio TOA. All the relevant time conversions are done by PINT, by using the parameter file created with the procedure in Section III. The phase of each photon can be calculated by properly converting the event arrival times from mission-elapsed seconds to MJD and using the TimingModel.phase method. Observations spanning multiple JBE solutions (e.g. from Fermi, IXPE) were split in the relevant time ranges. In very high-throughput observations with more than 100M photons (e.g. frequently from NICER) we calculated a TOA every $\sim$50M photons.
III.2 Deadtime correction of pulse profiles
@@
Figure 1: Example of deadtime correction of a NuSTAR pulse profile (ObsID 10002001009), using the method described in Section III.2. The blue line shows the exposure correction, the red line the raw profile and the black line the deadtime-corrected profile.
@@
For missions affected by dead time, like NuSTAR, we corrected the pulse profiles using the PRIOR column in the event list, which gives the live time before the current event, largely following the procedure used by Madsen et al. (2015). In short, the live time before each event was used to estimate the effective exposure in each bin of the pulse profile. For example, let us consider a photon arrived at time $t_{0}$, with a value in the PRIOR column of 1 ms. This means that the live time before this photon started at $t_{p}=t_{0}-1\,\mathrm{ms}$. Let us call $\phi_{0}$ the phase of the pulse profile corresponding to $t_{0}$, and $\phi_{p}$ the phase corresponding to $t_{p}$. Our procedure will add 1 to all phase bins between $\phi_{p}$ and $\phi_{0}$, properly accounting for fractional bin coverage and phase wraps at the start and end of each pulse profile. In the end, the exposure will be normalized to a maximum of 1, and the pulse profile divided by this effective exposure (so that more deadtime-affected pulse bins will be amplified). This procedure is made available in the publically available code pulse_deadtime_fix (Bachetti, 2025b). The result of this procedure is shown in Figure 1.
@@
between 12 and 24 counts per second to each bin of the folded profile, following Pease and Donnelly (1998).
III.3 Residual calculation
@@
Figure 2: Folded profile from NICER (obsid 101301013) modeled with the combination of symmetric and asymmetric Lorentzians described in Section III.3.
@@
Once we have the folded profile, as we saw in Section III.1, the phase 0 of the profile should coincide with the main radio peak. As is very well known from the literature (Rots et al., 2004), the peak in the X-rays will not coincide with the radio peak, and we need a way to measure this delay, that can then be used to compare the performance of different high-energy missions.
@@
There are mainly two methods used in the literature to calculate the phase of a pulsar: the alignment of the folded profile with a template (e.g. FFTFIT, Taylor 1992) or fitting a model to the pulse profile (e.g. Nelson et al., 1970). The first method is the most stable in ideal conditions, when the template for the pulse is correct, bringing precise measurements even in noisy data. However, if the template is not correct, as can happen if the response of the instrument used to calculate the template is different from the instrument used to obtain the folded data, or worse, if the folding is not perfect due to issues with the timing systems, this method can give unrealistically small errors and/or align to the wrong peaks. In the case of the Crab pulsar, where low signal-to-noise is seldom a problem, the second method yields meaningful results both in “good” datasets without requiring an exact template, and in datasets affected by instrumental issues, where the template adapts to an imperfect profile measuring the peak position and providing a reasonable error bar.
@@
For this work, we chose to model each peak as a combination of a symmetric Lorentzian curve plus an asymmetric Lorentzian curve. The phases of the two Lorentzians composing each peak are linked (not more than a difference of 0.02 in phase). All other parameters are left free to vary. The folded profile changes considerably at different energies, until the secondary peak becomes comparable to the first peak making the identification less obvious. However, the secondary peak maintains a distance of $\sim 0.4$ in phase from the primary, so that the two peaks can easily be identified by their phase distance. The error on the arrival time can be calculated from the error of the centroid of the sharpest Lorentzian. We verified that the scatter of TOA residuals is indeed comparable to the error bars estimated in this way. The fit quality can be seen in Figure 2.
III.4 Calculation of local pulsar frequency
We also calculated the local frequency of the pulsar by applying the quasi-fast folding algorithm (Bachetti et al., 2020), as implemented in Stingray (Huppenkothen et al., 2019), to each observation, searching a small frequency-frequency derivative plane around the nearest radio solution. We can use this information in special cases near a glitch or when the JBE is missing, and we plan to exploit it in future versions for the determination of X-ray-only ephemerides of the Crab untied to the JBE.
III.5 Validation
@@
Figure 3: Results of the validation procedure for the model refitting described in Section III.5. We fit a DE430 model to DE200-simulated TOAs, then simulated TOAs with the new model and calculated the residuals to the original DE200 model. For each instance of the simulation, we plot the mean residual in the upper panel and the rms of the residuals in the lower panel. The procedure works remarkably well, with residuals and rms comparable or lower to the injected error bars in the simulated data (0.1 $\rm\mu s$).
@@
Figure 4: End-to-end validation of the procedure in Section III using NICER datasets. (Left) Difference between the TOA residuals (labeled $\rm t_{eph}$) calculated from datasets barycentered using the DE200 ($\rm t_{200}$), DE430 ($\rm t_{430}$), and DE440 ($\rm t_{440}$) JPL ephemerides. (Left) a histogram of the corresponding cases. We did not consider observations where the TOA uncertainty was $15\,\rm\mu s$. The uncertainty of these TOAs was typically $\sim 5\,\rm\mu s$, and residual differences are always well inside error bars.
@@
Starting from a model of the Crab pulsar defined in the DE200 ephemeris, we simulated one month of geocentered TOAs as in Section III, we fit a DE430 model to them using a random position in the ICRS frame around the JBE position, simulated geocentered TOAs with the new model, and calculated the residuals to the initial model. Results are shown in Figure 3. We found that, for the method to work with a precision of better than 5 $\rm\mu s$, quadruple precision [^10] Here we refer to the data type often called float128, despite being in practice an 80-bit floating point in common architectures and, therefore, not actual quadruple precision. floating points have to be available in the computer doing the analysis. For example, in recent ARM-based Mac machines, the scatter in Figure 3 increased by a factor $\sim$20, with strong patterns dependent on the pulse frequency and the day of year. The problem did not appear on Intel-based Macs, x86 Linux-powered machines, or when using the Rosetta emulator on ARM-based Macs (which is also capable of emulating long doubles). Also, we found that the fit with the two-derivative model was always adequate, yielding an rms scatter of residual comparable with the injected 0.1 $\rm\mu s$ errorbars, if the simulated position was <5<5″, which is typically the case in our observations, and still $\lesssim$1 $\rm\mu s$ for a position error <1<1′.
End-to-end validation
We proceeded to an end-to-end validation the procedure using NICER observations of the Crab pulsar, barycentered using the DE200, DE430, and DE440 ephemerides and folded using the timing models obtained with our method. We folded the profiles and applied the residual calculation described in Section III.3, and then compared the results. We find that the difference between the residuals using the DE200 and DE4xx ephemerides was always consistent with 0 (Figure 4). This demonstrates that the method described in this paper is able to calculate TOAs from a variety of JPL ephemerides and positions, and that the differences are comparable to the TOA error at this level, with a possible systematic shift by $\sim 2\,\rm\mu s$ that we are investigating. This level of precision is adequate for Crab studies, but can be improved by analyzing sharper pulsars. This is planned for future studies.
III.6 Pipeline